# U-ARM : Ultra low-cost general teleoperation interface for robot manipulation

### v1

@@ Abstract
We propose U-Arm, a low-cost and rapidly adaptable leader-follower teleoperation framework designed to interface with most of commercially available robotic arms. Our system supports teleoperation through three structurally distinct 3D-printed leader arms that share consistent control logic, enabling seamless compatibility with diverse commercial robot configurations. Compared with previous open-source leader-follower interfaces, we further optimized both the mechanical design and servo selection, achieving a bill of materials (BOM) cost of only $50.5 for the 6-DoF leader arm and $56.8 for the 7-DoF version. To enhance usability, we mitigate the common challenge in controlling redundant degrees of freedom by mechanical and control optimizations. Experimental results demonstrate that U-Arm achieves 39% higher data collection efficiency and comparable task success rates across multiple manipulation scenarios compared with Joycon, another low-cost teleoperation interface. We have open-sourced all CAD models of three configs and also provided simulation support for validating teleoperation workflows. We also open-sourced real-world manipulation data collected with U-Arm. The project website is github.com/MINT-SJTU/LeRobot-Anything-U-Arm.
@@ 1 Introduction
Collecting large-scale high-quality manipulation data for dual-arm robots has long been a bottleneck in policy learning. Compared to simulation or human data, real-world robot data is the most directly applicable for training robust policies Bjorck et al. (2025). Among the various data collection methods, currently human demonstration remains the primary approach for acquiring such real-world data.
@@
In this paper, we present an ultra-low-cost, user-friendly, and broadly compatible leader-follower teleoperation system that enables researchers to rapidly build customized data collection pipelines for their own robotic platforms. Fig 1 shows our system overview. Specifically, we design three mechanically distinct leader arm configurations—two 6-DoF variants and one 7-DoF variant. Considering the structural constraints common to commercial robotic arms—namely, the need for satisfying Pieper’s criterion in 6-DoF systems and anthropomorphic joint arrangements in 7-DoF systems—we observe that, although link lengths vary across platforms, the orientation and sequencing of joints generally follow standardized conventions.
@@ 2 Related Work
Low Cost Teleoperation System for Bimanual Robot. Various teleoperation interfaces for bimanual robot manipulation have been developed in recent years. Common devices include VR headsets Cheng et al. (2024), Xiong et al. (2025), game controllers with joysticks, Spacemouse Liu et al. (2022), multi-camera mocap systems Handa et al. (2020), Qin et al. (2023), Song et al. (2020) , IMU-based controllers Wu et al. (2019), Laghi et al. (2018), and leader-follower teleoperation systems Wu et al. (2024), Zhao et al. (2023). Most of these devices control the robot in end-effector space, which are then converted into joint commands via inverse kinematics(IK). In contrast, leader-follower teleoperation systems directly map the joint configurations of a leader arm to the follower robot, thereby avoiding singularity problems during motion and physically constraining the operator’s movements to ensure that the collected demonstrations are robot-executable.
@@
However, teleoperation systems must balance cost, usability, and control accuracy. In Table 1, we summarize several representative teleoperation setups, comparing their hardware cost and qualitatively evaluating their usability for robot demonstration tasks.
| Device |
| ALOHA Zhao et al. (2023) | $24,000 | ✓ | ✓ | ✗ | ✓ | ||||||||
| DexPilot Handa et al. (2020) | $1700 | ✓ | ✓ | ✗ | ✗ | ||||||||
| VR Headset (Meta Quest 3) Meta (2025) | $500 | ✗ | ✓ | ✗ | ✗ | ||||||||
| Space Mouse 3Dconnexion (2025) | $220 | ✓ | ✗ | ✗ | ✓ | ||||||||
| GELLO Wu et al. (2024) | $270 | ✓ | ✓ | ✗ | ✓ | ||||||||
| Game Controller (Joycon) Box2AI (2025) | $20 | ✓ | ✗ | ✓ | ✓ | ||||||||
| U-Arm | $50.5 | ✓ | ✓ | ✓ | ✓ |
@@
High-cost leader-follower teleoperation systems such as ALOHA benefit from advanced motor drivers and control algorithms, which enable gravity compensation. This significantly improves the overall usability of the system and alleviates the challenges associated with controlling redundant degrees of freedom. However, the prohibitive cost of such systems hinders their scalability and broader adoption in the research community. GELLO can be seen as a low-cost alternative to ALOHA, aiming to support teleoperation across various commercial robot configurations. While game controller seems to be a sufficient cheap interface for simple manipulation tasks, we will demonstrate in the experiment section that it is less efficient than our device in actual use.
@@ 3.1 Hardware Design
Motivations. Most commercially available 6-DoF and 7-DoF robotic arms adopt one of only 3 standardized joint sequencing patterns, as illustrated in Fig. 2. We summarize in Table 2 a representative set of commercial robotic arms that are compatible with each of the three configurations. This design consistency arises from practical considerations such as satisfying Pieper’s criterion, optimizing the dexterous workspace, and achieving anthropomorphic arm structures. While the exact link lengths of different arms may vary, their joint type and ordering typically follow one of these three common configurations.
@@
| Config-2 | Dobot CR5, UR5, ARX R5∗, AUBO i5, JAKA Zu7 |
@@
Mechanical Design. Corresponding to the configs presented above, the overview of mechanical design of the U-Arm is shown in Fig. 3. All parts are printed using PLA in our experiments. Considering PLA’s strength, all components of the U-Arm have a minimum wall thickness of 4mm in order to ensure durability.
@@
Motor Modification In leader-follower teleoperation systems, joint resistance is a key factor that significantly affects the user experience. When the joint resistance of the leader arm is too high, operators find it difficult to perform smooth and continuous motions. On the other hand, if the resistance is too low, although the arm is easy to move intuitively, it may exhibit undesired passive motions in certain scenarios—particularly in Config-1 and Config-3.
@@ 3.2 Algorithm Design
Servo Encoder Adjustment and Calibration This work uses joint angle mapping for teleoperation. The Zhongling servos used in our system have an encoder range of 0-270°, and exceeding this range can lead to unpredictable behavior. Since the gearboxes have been removed, the servos cannot correct their position through command feedback. Therefore, before installation, we manually adjusted the servos to a neutral position around 135°, ensuring that the leader arm does not exceed the encoder range during normal operation.
Calibration and Filtering Even with identical joint configurations, different robotic arms exhibit significant structural differences, making it impossible to specify a unified initial position for all systems. Thus, during each teleoperation process, we initialize the leader arm near the follower arm’s pre-defined initial pose. At the start of each demonstration, the follower robot firstly moves to its initial pose, then the leader arm module is initialized and taking control. During operation,the joint angles command sent by the leader-arm are filtered and interpolated to account for slight disturbances in the encoder readings, ensuring smooth and accurate motion despite mechanical variations between systems.
@@ 4.1 Simulation Environment Adaption
We adapt U-Arm to ManiSkill Tao et al. (2025), which is built upon SAPIEN Xiang et al. (2020). This environment provides a safe setting for users to test teleoperation with U-Arm. Before operating the real robot, users can verify potential issues like errors in joint angle mapping. Additionally, users can also collect several demonstrations in this setting for utilizing emerging data scaling method like MimicGen Mandlekar et al. (2023).
We provide simulation examples for controlling 7 different robot arms at our website, including Arx-x5, Xarm, SO100, Panda. Some examples are shown in Fig. 4.
@@ 4.3 Real-World Experiment
Observations. Across all tasks, the U-Arm demonstrated a 39% reduction in operation time compared to the Joycon, without a substantial decrease in success rate. This improvement can be attributed to its more intuitive control scheme. In particular, for tasks involving richer variations in motion (e.g., grasping a Fanta and placing it on a shelf), the one-to-one mapping between the leader and follower arms in the U-Arm enabled smoother execution.
@@ 5 Conclusion
In this work, we present U-Arm, an open-source, ultra-low-cost leader arm system designed for teleoperation. By providing three distinct mechanical configurations, U-Arm can be adapted to most commercial robotic arms on the market. To address the control challenges introduced by redundant degrees of freedom, we modified the servos and applied an angle-mapping algorithm with filtering and calibration. These improvements enable for more stable and intuitive teleoperation. Experimental results demonstrate that U-Arm achieves 39% higher data collection efficiency and comparable task success rates across multiple manipulation scenarios compared with Joycon, another low-cost teleoperation interface.

### v2

@@ Abstract
We propose U-Arm, a low-cost and rapidly adaptable leader-follower teleoperation framework designed to interface with most of commercially available robotic arms. Our system supports teleoperation through three structurally distinct 3D-printed leader arms that share consistent control logic, enabling seamless compatibility with diverse commercial robot configurations. Compared with previous open-source leader-follower interfaces, we further optimized both the mechanical design and servo selection, achieving a bill of materials (BOM) cost of only $50.5 for the 6-DoF leader arm and $56.8 for the 7-DoF version. To enhance usability, we mitigate the common challenge in controlling redundant degrees of freedom by mechanical and control optimizations. Experimental results demonstrate that U-Arm achieveshttps://www.overleaf.com/project/687f198ea58dc9686630d322 3https://mp.weixin.qq.com/s/xV0HSDeuo9UH1N46nZNGUA9% higher data collection efficiency and comparable task success rates across multiple manipulation scenarios compared with Joycon, another low-cost teleoperation interface. We have open-sourced all CAD models of three configs and also provided simulation support for validating teleoperation workflows. We also open-sourced real-world manipulation data collected with U-Arm. The project website is github.com/MINT-SJTU/LeRobot-Anything-U-Arm.
@@ 1 Introduction
Collecting large-scale high-quality manipulation data for dual-arm robots has long been a bottleneck in policy learning. Especially, for vision-language-action models (VLAs) Black et al. (2024); Kim et al. (2024); Bjorck et al. (2025); Yang et al. (2025); Zhen et al. (2024); Liu et al. (2025a), the preliminary scaling law has been verified, and realizing the general Embodied AI requires the collection of massive amounts of real-world data. Compared to simulation or human data, real-world robot data is the most directly applicable for training robust policies Bjorck et al. (2025). Among the various data collection methods, currently human demonstration remains the primary approach for acquiring such real-world data.
@@
In this paper, we present an ultra-low-cost, user-friendly, and broadly compatible leader-follower teleoperation system that enables researchers to rapidly build customized data collection pipelines for their own robotic platforms. Fig.1 shows our system overview. Specifically, we design three mechanically distinct leader arm configurations—two 6-DoF variants and one 7-DoF variant. Considering the structural constraints common to commercial robotic arms—namely, the need for satisfying Pieper’s criterion in 6-DoF systems and anthropomorphic joint arrangements in 7-DoF systems—we observe that, although link lengths vary across platforms, the orientation and sequencing of joints generally follow standardized conventions.
@@ 2 Related Work
2.1 Manipulation Data Collection
In the field of robotic manipulation, researchers are currently focused on the data-driven paradigm of imitation learning. As a result, various approaches to data collection have been explored to gather robot datasets. Learning from human videos has recently emerged as a promising direction. Works such as GR00TBjorck et al. (2025) and EgoVLAYang et al. (2025) leverage large-scale human-centric video datasets for policy training. At the same time, works like EgoMimicKareer et al. (2025), EgozeroLiu et al. (2025b), and DexcapWang et al. (2024) propose more robot-oriented methods for collecting such human-centric video data.
The advantage of this approach is that it enables knowledge transfer from the vast amount of human demonstration videos available online, and it holds the potential to scale up further with the development of more integrated hardware. However, this comes at the cost of introducing a significant human-robot embodiment gap. Consequently, such data can either be used primarily for pretraining large VLA modelsBjorck et al. (2025) or require heavy reliance on post-processing of videos and trajectory data to map human demonstrations into distributions more closely aligned with actual robot executions.
Another approach considered for scaling up robot data is leveraging simulation platforms such as Isaac SimNVIDIA Corporation (2025), MuJoCoTodorov et al. (2012), and SAPIENXiang et al. (2020). The most direct way to leverage simulation platforms for data augmentation is through scripted policies and domain randomizationMandlekar et al. (2023); Jiang et al. (2025), which collect task demonstrations across different scenes and object placements. Some works have also utilized the capabilities of generative models to support the generation of new tasks and scenes, optimizing policies in the processWang et al. (2023); Li et al. (2024); Hua et al. (2024).
The greatest advantage of simulation environments is that researchers can easily access all ground-truth information within the scene for policy training and planning. Current rendering technologies have made it possible for simulation environments to achieve photo-realistic fidelityNVIDIA Corporation (2025), closely resembling real-world scenarios. However, for high-dynamic, non-rigid, and contact-rich tasks with complex mechanical properties, simulators still face challenges in accurately simulating and executing policies, leaving a significant sim2real gap between simulated and real-world data.
2.2 Teleoperation System for Bimanual Robot
Embodiment and sim2real gap reflect the importance of collecting demonstrations directly from real-world through teleoperation. A key advantage of real-world teleoperation data lies in the separation between control and recording: teleoperation devices are typically used only for control, while the data for learning is collected directly from the robot executing the task. This ensures that the recorded demonstrations share the closest distribution with deployment. However, real-world teleoperation inevitably involves human operators, making it the most costly and time-consuming method compared with simulation or vision-based data sources. Consequently, research on teleoperation devices mainly focused on reducing the cost of data collection, improving operator efficiency, and designing more user-friendly interfaces, all while ensuring high quality of the collected demonstrations.
@@ 2.2 Teleoperation System for Bimanual Robot
In recent years, a variety of teleoperation interfaces for bimanual robot manipulation have been developed to facilitate such data collection. Common devices include VR headsets Cheng et al. (2024); Xiong et al. (2025), game controllers with joysticks, Spacemouse Liu et al. (2022), multi-camera mocap systems Handa et al. (2020); Qin et al. (2023); Song et al. (2020), IMU-based controllers Wu et al. (2019); Laghi et al. (2018), and leader-follower teleoperation systems Wu et al. (2024); Zhao et al. (2023).
Most of these devices control the robot in end-effector space, which is then mapped to joint configurations through inverse kinematics (IK). In contrast, leader-follower teleoperation systems directly map the joint configurations of a leader arm to a follower robot, thereby avoiding singularity problems during motion and physically constraining the operator’s movements to ensure that collected demonstrations remain robot-executable.
@@
Nevertheless, teleoperation systems must strike a balance between cost, usability, and control accuracy. In Table 1, we summarize several representative teleoperation setups, comparing their hardware cost and qualitatively evaluating their usability for robot demonstration tasks.
| Device | Price (USD) | Remarks |
| VR Headset (Meta Quest 3) Meta (2025) | $500 | May cause motion sickness |
| Space Mouse 3Dconnexion (2025) | $220 | Hard to operate bimanually |
| GELLO Wu et al. (2024) | $270 | Uses Dynamixel motors, $24 each |
| Game Controller (Joycon) Box2AI (2025) | $20 | Difficult for dexterous manipulation |
| Ours | $50.5 | Sufficient for bimanual manipulation |
| ALOHA Zhao et al. (2023) | $24,000 | Identical leader-follower hardware design |
| DexPilot Handa et al. (2020) | $1700 | Faces occlusion problems |
@@
High-cost leader-follower teleoperation systems such as ALOHA benefit from advanced motor drivers and control algorithms, which enable gravity compensation. This significantly improves the overall usability of the system and alleviates the challenges associated with controlling redundant degrees of freedom. However, the prohibitive cost of such systems hinders their scalability and broader adoption in the research community.
2.3 Low-cost Teleoperation System
GELLO Wu et al. (2024) can be considered as a low-cost alternative to ALOHA, aiming to support teleoperation across various commercial robot configurations. While game controller seems to be a sufficient cheap interface for simple manipulation tasks, we will demonstrate in the experiment section that it is less efficient than our device in actual use.
@@ 3.1 Hardware Design
Motivations
Most commercially available 6-DoF and 7-DoF robotic arms adopt one of only 3 standardized joint sequencing patterns, as illustrated in Fig. 2. We summarize in Table 2 a representative set of commercial robotic arms that are compatible with each of the three configurations. This design consistency arises from practical considerations such as satisfying Pieper’s criterion, optimizing the dexterous workspace, and achieving anthropomorphic arm structures. While the exact link lengths of different arms may vary, their joint type and ordering typically follow one of these three common configurations.
@@
| Config-2 | Dobot CR5, UR5, ARX RS5*S, AUBO i5, JAKA Zu7 |
@@
Mechanical Design
Corresponding to the configs presented above, the overview of mechanical design of the U-Arm is shown in Fig. 3. All parts are printed using PLA in our experiments. Considering PLA’s strength, all components of the U-Arm have a minimum wall thickness of 4mm in order to ensure durability.
@@
Motor Modification
In leader-follower teleoperation systems, joint resistance is a key factor that significantly affects the user experience. When the joint resistance of the leader arm is too high, operators find it difficult to perform smooth and continuous motions. On the other hand, if the resistance is too low, although the arm is easy to move intuitively, it may exhibit undesired passive motions in certain scenarios—particularly in Config-1 and Config-3.
@@ 3.2 Algorithm Design
Servo Encoder Adjustment and Calibration
This work uses joint angle mapping for teleoperation. The Zhongling servos used in our system have an encoder range of 0-270°, and exceeding this range can lead to unpredictable behavior. Since the gearboxes have been removed, the servos cannot correct their position through command feedback. Therefore, before installation, we manually adjusted the servos to a neutral position around 135°, ensuring that the leader arm does not exceed the encoder range during normal operation.
Calibration and Filtering
Even with identical joint configurations, different robotic arms exhibit significant structural differences, making it impossible to specify a unified initial position for all systems. Thus, during each teleoperation process, we initialize the leader arm near the follower arm’s pre-defined initial pose. At the start of each demonstration, the follower robot firstly moves to its initial pose, then the leader arm module is initialized and taking control. During operation,the joint angles command sent by the leader-arm are filtered and interpolated to account for slight disturbances in the encoder readings, ensuring smooth and accurate motion despite mechanical variations between systems.
@@ 4.1 Simulation Environment Adaption
We adapt U-Arm to ManiSkill Tao et al. (2025), which is built upon SAPIEN Xiang et al. (2020). This environment provides a safe setting for users to test teleoperation with U-Arm in different simulation scenes. Before operating the real robot, users can verify potential issues like errors in joint angle mapping and get easy access to teleoperation without the need of a real-world robot. Additionally, users can also collect several demonstrations in this setting for utilizing emerging data scaling method like MimicGen Mandlekar et al. (2023).
We set up the simulation environment in an indoor setting with various desktop items, where we test the capacity of the U-Arm to teleoperate robot arms to accomplish basic tabletop manipulation tasks and prove the cross-model generalization of U-Arm.
We provide simulation examples for controlling 7 different robot arms at our website, including ARX-X5, Xarm6, SO-100, Panda. Some examples are shown in Fig. 4.
@@ 4.3 Real-world Experiment
Observations
Across all tasks, the U-Arm demonstrated a 39% reduction in operation time compared to the Joycon, without a substantial decrease in success rate. This improvement can be attributed to its more intuitive control scheme. In particular, for tasks involving richer variations in motion (e.g., grasping a Fanta and placing it on a shelf), the one-to-one mapping between the leader and follower arms in the U-Arm enabled smoother execution.
@@
Trajectory Smoothness
As illustrated in Fig. 6, the end-effector trajectories obtained from the U-Arm interface follow smoother and more continuous paths. In contrast, trajectories collected with the Joycon controller typically consist of piecewise, axis-aligned movements, reflecting the discrete nature of joystick inputs. This observation suggests that leader-follower teleoperation system induces a trajectory distribution that is closer to natural human manipulation. Such a property is particularly desirable if one aims to co-train policies with other data sources (e.g., ego-centric human demonstrationsWang et al. (2024)), since the data distribution collected via the U-Arm better aligns with the statistics of human motion.
Analysis of Operator Proficiency
It must be admitted that though some work are exploring shared autonomy in bimanual manipulation tasksLuo et al. (2025), currently teleoperation system still highly rely on human labor and experience. To further investigate the effect of operator proficiency, we compare success rates across five trial ranges (1-10, 11-20, 21-30, 31-40, 41-50), as shown in Fig. 7. We observe that U-Arm shows a clear improvement after a short adaptation phase: while its initial success rate is slightly lower, it steadily increases and stabilizes around 80% after 10-20 trials. In contrast, the Joycon exhibits relatively high success rates from the beginning, indicating that it is easier for operators to quickly adapt. Nevertheless, the U-Arm achieves comparable steady-state performance after a short learning period, demonstrating that its intuitive control scheme remains effective once operators become familiar with it.
@@ 5 Conclusion
In this work, we present U-Arm, an open-source, ultra-low-cost leader arm system designed for teleoperation. By providing three distinct mechanical configurations, U-Arm can be adapted to most commercial robotic arms on the market. To address the control challenges introduced by redundant degrees of freedom, we modified the servos and applied an angle-mapping algorithm with filtering and calibration. These improvements enable for more stable and intuitive teleoperation.
Experimental results demonstrate that U-Arm achieves 39% higher data collection efficiency and comparable task success rates across multiple manipulation scenarios compared with Joycon, another low-cost teleoperation interface. Meanwhile, as a leader-follower teleoperation system, U-Arm performs trajectories that more closely reflect natural human motions compared with Joycon. This property makes U-Arm particularly advantageous when co-training policies with other human-centric data sources.